OpenStudy (lilsis76):

Use the formula cos(s+t)=cos s cos t - sin s sin t: to compute the exact value of cos 75 degrees. all i know is that 75 degrees is between 90 deg and 60 degrees on a unit circle.

zepdrix (zepdrix):

Hmmmmm we have special angles for 30, 45, 60, 90... do any of those 2 add up to 75? Cause that would really help us out in this problem.

OpenStudy (lilsis76):

well... okay we have the 30 and 45 angles

zepdrix (zepdrix):

Hmm good good :) I think we can work with that. So we'll rewrite cos(75) as cos(45+30). After we apply the sum formula, we should be able to plug in a BUNCH of special angles and HOPEFULLY we'll be able to simplify it :3 hehe

zepdrix (zepdrix):

So use your formula you posted at the top, what will your setup look like? :D

OpenStudy (lilsis76):

okay.

OpenStudy (lilsis76):

the 45 will be the s, and t will be tht 30 so... cos(s+t)=cos s cos t - sin s sin t: cos(45+30)= cos 45 cos 30 - sin 45 sin 30. yes??

zepdrix (zepdrix):

Mmmmmmmmmmmmmm yah looks good. Now it's special angle time! :OOO

OpenStudy (lilsis76):

what does that mean?

zepdrix (zepdrix):

Let's look at the first ..thing.. cos(45), what is the cosine of 45 degrees? It's a special angle that you're suppose to remember.. (from the unit circle) :X

OpenStudy (lilsis76):

OH....so the points are the special angles??? I thought that meant another formula to remember. um... 45 is at squ rt. 2 / squ rt. 2 / 2 cos sin cos is squ.rt. 2 / 2

zepdrix (zepdrix):

so cos(45)=sqrt2/2 and cos(30)? :D Go through all of them and do yer magic!

OpenStudy (lilsis76):

what?! the 30 ? um... hold on

OpenStudy (lilsis76):

squ rt 3 /2

OpenStudy (lilsis76):

okay hold on lol

zepdrix (zepdrix):

so our first term cos(45)cos(30) becomes.. (sqrt2/2)(sqrt3/2) Which we should be able to simplify a little bit. hmm

OpenStudy (lilsis76):

okay tis.. (sqrt2 /2)( sqrt3/2) - (sqrt 2/2)(1/2)

zepdrix (zepdrix):

k good good c: Now simplyyyyyyyyyy \:D/

OpenStudy (lilsis76):

uh...hold on let me try doin it um..

OpenStudy (lilsis76):

i get sqrt3 /2 - 1/2

zepdrix (zepdrix):

 Hmmmm this is what I'm coming up with D: